OpenStudy (anonymous):

what is the domain of ln (2x^2 -4)-ln(x+1)-ln(x-1)

OpenStudy (dan815):

fine the places there the inner brackets equal 0 because ln (0) = DNE

OpenStudy (zzr0ck3r):

domain of ln(x) is x>0 so you need the intersection of these three 2x^2-4>0 x+1>0 and x-1>0

OpenStudy (dan815):

the inner brackets of each expression must be > than 0

OpenStudy (dan815):

and then do what he says ^ lol

OpenStudy (dan815):

just make sure u give meee the MEDAL in the end! :D

OpenStudy (anonymous):

i worked out the x values of each ln what is the next step the - between the ln confuses me

OpenStudy (dan815):

solve the 3 brackets where x>0 and see which of those x values will be valid for all of them

OpenStudy (anonymous):

so if x = +/- 2 and +/- 1 how do we end the answer and what is the domain?

OpenStudy (dan815):

basically x must be greater than > sqrt2

OpenStudy (zzr0ck3r):

x+1>0 and x-1>0 implies x>1 and x>-1 and 2x^2-4>0 imples x<-sqt(2) or x>sqrt(2) so x>1 and x>-1 and (x>sqrt(2) or x<-sqr(2) so x>sqrt(2) is the domain

OpenStudy (anonymous):

(-OO; 2) ? is the domain?

OpenStudy (zzr0ck3r):

\[(\sqrt(2),infinity)\]

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

because x>1 and x>-1 so x cant be <-1

OpenStudy (anonymous):

Well, \(\displaystyle\ln x= ln|x|+ i2\arctan\left(\frac{x-\overline x}{2i\sqrt{\left(\dfrac{x+\overline x}{2}\right)^2+\left(\dfrac{x-\overline x}{2i}\right)^2}+\dfrac{x+\overline x}{2}}\right)\) even when x<0.